# SF-V: Single Forward Video Generation Model

### v1

@@ 1 Introduction
Video generation is experiencing unprecedented advancements by leveraging large-scale denoising diffusion probabilistic models [1, 2] to create photo-realistic frames with natural and consistent motion [3, 4], revolutionizing various fields, such as entertainment and digital content creation [5].
@@ 3.2 Latent Adversarial Training for Video Diffusion Model
$\hat{x}_{0}(x_{t};\sigma_{t},\mathbf{c})=c_{skip}(\sigma_{t})x_{0}+c_{out}(\sigma_{t})\mathcal{G}_{\theta}(c_{in}(\sigma_{t})x_{t};c_{noise}(\sigma_{t}),\mathbf{c}).$ (6)
@@
Discussion. Our latent adversarial training framework is largely inspired by LADD [27]. Similar to LADD, we set $T_{g}=4$ in practice and utilize a pre-trained diffusion model as part of the discriminator. However, our approach has several key differences compared with LADD. First, we extend the image latent adversarial distillation framework to the video domain by incorporating spatial and temporal heads to achieve one-step generation for video diffusion models. The specifics of the spatial and temporal heads are discussed in Sec. 3.3. Additionally, we notice that, using the UNet encoder as the discriminator backbone, instead of the entire UNet as in LADD [27], is sufficient to achieve high-quality results. This smaller discriminator design is especially important for video diffusion model as it enables faster training and consumes less GPU memory. Second, based on the EDM-framework [32], we observe that sampling $t^{\prime}$ using a discretized lognormal distribution provides more stable adversarial training compared to the logit-normal distribution used in LADD [27]. Finally, unlike LADD [27], we utilize real video data instead of synthetic data for training and incorporate a reconstruction objective (i.e., Eq. (9)) to ensure more stable training.
@@ 3.3 Spatial Temporal Heads
Table 1: Comparison Results. We compare our method against SVD [12] and AnimateLCM [20] using different numbers of sampling steps. AnimateLCM indicates the usage of the officially provided $25$-frame model, with only the first $14$ frames considered for FVD calculation. We also report the latency of the denoising process for each setting, measured on a single NVIDIA A100 GPU.
@@ 4.2 Comparisons Results
Figure 5: Comparison between SVD [12], AnimateLCM [20], and Our Approach. We provide the synthesized videos (sampled frames) under various settings for different approaches. We use SVD to generate videos under $25$, $16$, and $8$ sampling steps, and AnimateLCM to synthesize videos under $4$ sampling steps. AnimateLCM generates blurry images when applied with few sampling steps. Our approach can accelerate the sampling speed by 22.9×22.9$\times$22.9 × compared with SVD while maintaining similar frame quality and motion consistency.
Quantitative Comparisons. We present a comprehensive evaluation of our method compared to the existing state-of-the-art approach, AnimateLCM [20], and SVD [12]. To conduct a fair comparison on the SVD model, we train the AnimateLCM on SVD using our video dataset. We follow the released code and instructions provided by AnimateLCM authors. Additionally, we include the officially released AnimateLCM-xt1.1 [20] by evaluating the first $14$ generated frames and denote the approach as AnimateLCM.
@@
As shown in Tab. 1, our method achieves comparable results to the base model using $16$ discrete sampling steps, resulting in approximately a 23×23$\times$23 × speedup. Our method also outperforms the $8$-steps sampling results for AnimateLCM and AnimateLCM, indicating a speedup of more than 6×6$\times$6 ×.
Qualitative Comparisons. We further provide qualitative comparisons across different approaches by using publicly available web images. Fig. 5 presents generation results from SVD [12] with $25$, $16$, and $8$ sampling steps, AnimateLCM [20] with $4$ sampling steps, and our method with $1$ sampling step. As can be seen, our method achieves results comparable to the sampling results of SVD using $16$ or $25$ denoising steps. We notice significant artifacts for videos synthesized by SVD when using $8$ denoising steps. Compared to AnimateLCM, our method produces frames of higher quality and better temporal consistency, with even fewer denoising steps, demonstrating the effectiveness of our proposed approach.
@@ 5 Discussion and Conclusion
Limitations. This work successfully achieves single sampling step for video diffusion models. However, under such setting, the temporal VAE decoder and the encoder for image conditioning take a considerable portion of the overall runtime. We leave the acceleration of these models as future work.

### v2

@@ 1 Introduction
Video generation is experiencing unprecedented advancements by leveraging large-scale denoising diffusion probabilistic models [1, 2] to create photo-realistic frames with natural and consistent motion [3, 4], revolutionizing various fields, such as entertainment and digital content creation [5, 6].
@@ 3.2 Latent Adversarial Training for Video Diffusion Model
$\hat{x}_{0}(x_{t};\sigma_{t},\mathbf{c})=c_{skip}(\sigma_{t})x_{t}+c_{out}(\sigma_{t})\mathcal{G}_{\theta}(c_{in}(\sigma_{t})x_{t};c_{noise}(\sigma_{t}),\mathbf{c}).$ (6)
@@
Discussion. Our latent adversarial training framework is largely inspired by LADD [28]. Similar to LADD, we set $T_{g}=4$ in practice and utilize a pre-trained diffusion model as part of the discriminator. However, our approach has several key differences compared with LADD [28]. First, we extend the image latent adversarial distillation framework to the video domain by incorporating spatial and temporal heads to achieve one-step generation for video diffusion models. The specifics of the spatial and temporal heads are discussed in Sec. 3.3. Second, based on the EDM-framework [33], we observe that sampling $t^{\prime}$ using a discretized lognormal distribution provides more stable adversarial training compared to the logit-normal distribution used in LADD [28]. Finally, unlike LADD [28], we utilize real video data instead of synthetic data for training and incorporate a reconstruction objective (i.e., Eq. (9)) to ensure more stable training.
@@ 3.3 Spatial Temporal Heads
Table 1: Comparison Results. We compare our method against SVD [13], AnimateLCM [21], UFOGen [25], and LADD [28] using different numbers of sampling steps. AnimateLCM indicates the usage of the officially provided $25$-frame model, with only the first $14$ frames considered for FVD calculation. indicates our implementations. We also report the latency of the denoising process for each setting, measured on a single NVIDIA A100 GPU.
@@ 4.2 Comparisons Results
Figure 5: Comparison between SVD [13], AnimateLCM [21], LADD [28], UFOGen [25], and Our Approach. We provide the synthesized videos (sampled frames) under various settings for different approaches. We use SVD to generate videos under $25$, $16$, and $8$ sampling steps, AnimateLCM to synthesize videos under $4$ sampling steps, LADD and UFOGen to generate videos under $1$ sampling step. AnimateLCM, LADD and UFOGen generates blurry frames with few-steps and single-step sampling. Our approach can accelerate the sampling speed by 22.9×22.9$\times$22.9 × compared with SVD while maintaining similar frame quality and motion consistency.
Quantitative Comparisons. We present a comprehensive evaluation of our method compared to the existing state-of-the-art approach, AnimateLCM [21], UFOGen [25], LADD [28], and SVD [13]. To conduct a fair comparison on the SVD model, we train the AnimateLCM, UFOGen, and LADD on SVD using our video dataset. We follow the released code and instructions provided by AnimateLCM authors. Additionally, we include the officially released AnimateLCM-xt1.1 [21] by evaluating the first $14$ generated frames and denote the approach as AnimateLCM. We try our best to implement LADD [28] and UFOGen [25] and denote respectively as LADD, and UFOGen. Note that simply re-using the discriminator from LADD [28] and UFOGen [25] leads to out-of-memory issue, since the computation in the video model is much larger than the image model. Here we replace the discriminator from LADD [28] and UFOGen [25] with the one proposed in our work.
@@
As shown in Tab. 1, our method achieves comparable results to the base model using $16$ discrete sampling steps, resulting in approximately a 23×23$\times$23 × speedup. Our method also outperforms the $8$-steps sampling results for AnimateLCM and AnimateLCM, indicating a speedup of more than 6×6$\times$6 ×. For single-step evaluation, our method performs much better than existing step-distillation methods [25, 28] built upon image-based-diffusion models.
Qualitative Comparisons. We further provide qualitative comparisons across different approaches by using publicly available web images. Fig. 5 presents generation results from SVD [13] with $25$, $16$, and $8$ sampling steps, AnimateLCM [21] with $4$ sampling steps, UFOGen [25], LADD [28], and our method with $1$ sampling step. As can be seen, our method achieves results comparable to the sampling results of SVD using $16$ or $25$ denoising steps. We notice significant artifacts for videos synthesized by SVD when using $8$ denoising steps. Compared to AnimateLCM [21],UFOGen [25], and LADD [28], our method produces frames of higher quality and better temporal consistency, with fewer or same denoising steps, demonstrating the effectiveness of our proposed approach.
@@ 5 Discussion and Conclusion
Figure 8: Limitations. We show that, for some conditional images, our model tends to generate a few unsatisfactory frames when complex motion might be required (Second Row). Similar artifacts can also be observed in frames generated from SVD by sampling at $25$-steps (First Row).
Limitations. We observe that when the given conditioning image indicates complex motion, e.g.running, our model tends to generate unsatisfactory results, e.g.blurry frames, as shown in Fig. 8. Such artifacts are introduced by the original SVD model, as can be observed in Fig. 8. We believe a better text-to-video model can solve such issue.
This work successfully achieves single sampling step for video diffusion models. However, under such setting, the temporal VAE decoder and the encoder for image conditioning take a considerable portion of the overall runtime. We leave the acceleration of these models as future work.